OpenStudy (anonymous):

Can someone help me factor this? x^6-y^6

OpenStudy (anonymous):

Use difference of squares, then sum and difference of cubes formulas.

OpenStudy (anonymous):

\[x^6−y^6=(x^3)^2−(y^3)^2=(x^3+y^3)(x^3−y^3) \]\[=(x+y)(x−y)(x^2−xy+y^2)(x^2+xy+y^2)\]